OpenStudy (anonymous):

how to solve (2y+3)^2=9

OpenStudy (anonymous):

First, square root both sides and you get 2y+3=3 Now subtract 3 on both sides 2y=0 Now divide by 2 y=0 is your final answer

OpenStudy (anonymous):

(2y+3)(2y+3) = 9 4y^2+6y+6y+9 = 9 4y^2+12y+9 = 9 4y^2+12 = 0 from there use quadratic formula or factoring

OpenStudy (anonymous):

* 4y^2 + 12y = 0

OpenStudy (aum):

\((2y+3)^2=9\) Take square root on both sides: \(2y + 3 = \pm 3\) 2y + 3 = 3 2y = 0 y = 0 2y + 3 = -3 2y = -6 y = -3 y = -3 and y = 0.

OpenStudy (anonymous):

thank y'all (:

OpenStudy (anonymous):

how to solve for \[6x^2+2x-1=0\]